# An Elliott intertwining approach to classifying actions of $\mathrm{C}^{*}$-tensor categories

### 29 Jan 2024 v2

@@ 3. The generalised cocycle category
In this section we introduce the category of $\mathcal{C}$-$\mathrm{C}^{*}$-algebras for which we will later perform intertwining arguments. Throughout the rest of this paper $\mathcal{C}$ is always assumed to be a $\mathrm{C}^{*}$-tensor category. Recall the definition of an action of $\mathcal{C}$ on a $\mathrm{C}^{*}$-algebra $A$ from Definition 2.1.
@@
[…]
@@
    A correspondence morphism from $(A,F,J)$ to $(B,G,I)$ is a pair $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}})$, where $\phi:A\to\mathcal{L}(E)$ is an $A$-$B$ correspondence and $\{\mathbbm{v}_{X}:\phi\circ F(X)\to G(X)\circ\phi\}_{X\in\mathcal{C}}$ is a natural family of $A$-$B$-bimodule maps such that $\mathbbm{v}_{X}$ is an isometry (not necessarily adjointable) for any $X\in\mathcal{C}$ [^6] By an isometry, we mean a map which preserves the norm. By the polarisation identity, it is equivalent to assume that it preserves the inner product. Moreover, for all $X,Y\in\mathcal{C}$ the following pentagon diagram commutes
    (3.1) {tikzcd}
@@
    A cocycle representation $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ is a correspondence morphism for which we further require that $\phi:A\to\mathcal{M}(B)$ is a $*{}^{*}$-homomorphism.
@@
    A cocycle morphism $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ is a cocycle representation for which we further require that $\phi:A\to B$ is a $*{}^{*}$-homomorphism.
@@
Example 3.3.
Suppose $(A,\alpha)$ and $(B,\beta)$ are actions of a group $\Gamma$ on $\mathrm{C}^{*}$-algebras $A$ and $B$. Consider them as actions of $\mathrm{Hilb}(\Gamma)$ as in Example 2.3. Let $(\phi,\mathbbm{v}):(A,\alpha)\to(B,\beta)$ be an extendible cocycle morphism as in Definition 3.1 with $\mathbbm{v}_{g}$ being adjointable for all $g\in\Gamma$. Fix $g\in\Gamma$ and let $f_{g}:=T_{g}^{\prime}\circ\mathbbm{v}_{\mathbb{C}_{g}}\circ S_{g}:{}_{\phi\circ\alpha_{g}}B\to{}_{\beta_{g}\circ\phi}B$, where $S_{g}:{}_{\phi\circ\alpha_{g}}B\to{}_{\alpha_{g}}A\boxtimes{}_{\phi}B$ is given by $S_{g}(b)=\lim\limits_{\lambda}\xi_{\lambda}\boxtimes b$ for any $b\in B$, where $\xi_{\lambda}$ is an approximate unit for $A$ and $T_{g}^{\prime}:{}_{\phi}B\boxtimes{}_{\beta_{g}}B\rightarrow{}_{\beta_{g}\circ\phi}B$ is given by $T_{g}^{\prime}(b\boxtimes c)=\beta_{g}(b)c$ for $b,c\in B$ [^8] Note that $S_{g}$ is well-defined by Proposition 1.7. Moreover, let $T_{g}:{}_{\alpha_{g}}A\boxtimes{}_{\phi}B\rightarrow{}_{\phi\circ\alpha_{g}}B$ be given by $T_{g}(a\boxtimes b)=\phi(a)b$ for any $a\in A$ and any $b\in B$.
@@
[…]
@@
Note that, unlike in [40, Definition 1.10], we do not require $\phi\in\mathrm{Corr}(A,B)$ to be extendible. In fact, this is precisely the reason why we consider the maps $\mathbbm{v}_{X}$ to be isometries (possibly non-adjointable) instead of unitaries. For example, the map $\mathbbm{v}_{1_{\mathcal{C}}}:A\boxtimes{}_{\phi}B\to{}_{\phi}B\boxtimes B\cong{}_{\phi}B$ is given by $\mathbbm{v}_{1_{\mathcal{C}}}(a\boxtimes b)=\phi(a)b$ for all $a\in A$ and $b\in B$. Therefore, it is not surjective unless $\phi$ is non-degenerate. Moreover, it might not be adjointable if $\phi$ is not extendible as seen in the proof of Proposition 1.7.
Furthermore, the morphisms of Definition 3.1 fit into the $\mathcal{C}$-equivariant $\mathrm{KK}$-theory developed in [1] (see [1, Example 3.3]). In [1] a correspondence morphism is instead called a $\mathcal{C}$-Hilbert $A$-$B$-bimodule and a cocycle morphism is called a cocycle-$\mathcal{C}$-***-homomorphism.
[…]
@@
By combining the pentagon diagrams for $\mathbbm{v}$ and $\mathbbm{w}$, one obtains the following lemma.
Lemma 3.5.
With the notation above, $(\psi\circ\phi,\mathbbm{w}\circ\mathbbm{v}):(A,F,J)\to(C,H,K)$ is a correspondence morphism which denotes the composition of $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$.
@@
[…]
Here $S_{X}(x\boxtimes c)=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}\boxtimes c$ and $T(b\boxtimes c)=\psi(b)c$ for any $X\in\mathcal{C}$, $x\in F(X)$, $b\in B$, and $c\in C$, with $\eta_{\lambda}$ being an approximate unit for $B$ [^10] We have not shown that $S_{X}$ is well-defined at this point.
@@
By the Cauchy-Schwarz inequality and since $(\eta_{\lambda}-\eta_{\mu})\phi(\langle x,x\rangle_{A})^{1/2}$ converges to 00, it is readily seen that $\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c,x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c\rangle_{C}$ converges to 00, so $S_{X}$ is well-defined. Moreover, commutation with the left and right actions are immediate.
@@
Therefore, by linearity and continuity of both $S_{X}$ and $T_{X}$, it follows that $S_{X}$ is surjective and $T_{X}$ is injective. It now suffices to show that $T_{X}$ is surjective. This will imply that $S_{X}$ is invertible with $T_{X}$ being the inverse. Let $x\in F(X)$ and $c\in C$. Since $F(X)$ is a non-degenerate $A$-bimodule, recall from Lemma 1.11 that the map $R_{X}^{-1}:F(X)\boxtimes A\to F(X)$ given by $R_{X}^{-1}(x\boxtimes a)=x\lhd a$ is a bimodule isomorphism. Then, let $y\in F(X)$ and $a\in A$ such that $y\lhd a=x$. A straightforward calculation shows that $T_{X}(y\boxtimes\phi(a)\boxtimes c)=x\boxtimes c$, so $T_{X}$ is surjective. Hence, $S_{X}=T_{X}^{-1}$ is an isomorphism for any $X\in\mathcal{C}$.
@@
Note that the proof of Lemma 3.7 also shows that the map $S_{X}$ does not depend on the choice of approximate unit.
@@
Proof.
Note that $\psi\circ\phi:A\to C$ is a $*{}^{*}$-homomorphism. Then, $(\mathbbm{w}*\mathbbm{v})_{X}$ is an isometry as it is a composition of isometries, and it is also an $A$-$C$-bimodule map as it is a composition of bimodule maps. Moreover, naturality of $\mathbbm{w}*\mathbbm{v}$ follows from naturality of $\mathbbm{w}\circ\mathbbm{v}$. Let us now check that the family of maps $\{(\mathbbm{w}*\mathbbm{v})_{X}\}$ satisfies the diagram
(3.5) {tikzcd}
for all $X,Y\in\mathcal{C}$.
Fix $X,Y\in\mathcal{C}$. By Definition 3.6, composing the two rightmost arrows in (3.5) yields that
(3.6) $(\mathbbm{w}*\mathbbm{v})_{X\otimes Y}\circ(J_{X,Y}\boxtimes\mathrm{id}_{C})$
$=(T\boxtimes\mathrm{id}_{H(X\otimes Y)})\circ(\mathbbm{w}\circ\mathbbm{v})_{X\otimes Y}\circ S_{X\otimes Y}\circ(J_{X,Y}\boxtimes\mathrm{id}_{C})$
(3.7) $=(T\boxtimes\mathrm{id}_{H(X\otimes Y)})\circ(\mathbbm{w}\circ\mathbbm{v})_{X\otimes Y}\circ(J_{X,Y}\boxtimes\mathrm{id}_{B}\boxtimes\mathrm{id}_{C})\circ(\mathrm{id}_{F(Y)}\boxtimes S_{X}).$
Then, by Lemma 3.5,
(3.8) $(\mathbbm{w}\circ\mathbbm{v})_{X\otimes Y}\circ(J_{X,Y}\boxtimes\mathrm{id}_{B}\boxtimes\mathrm{id}_{C})$
(3.9) $=(\mathrm{id}_{B}\boxtimes\mathrm{id}_{C}\boxtimes K_{X,Y})\circ\big{(}(\mathbbm{w}\circ\mathbbm{v})_{Y}\boxtimes\mathrm{id}_{H(X)}\big{)}\circ\big{(}\mathrm{id}_{F(Y)}\boxtimes(\mathbbm{w}\circ\mathbbm{v})_{X}\big{)}.$
Substituting (3.8) into (3.6), and using Definition 3.6, it follows that the diagram in (3.5) commutes.
It remains to check that $(\mathbbm{w}*\mathbbm{v})_{1_{\mathcal{C}}}:A\boxtimes{}_{\psi\circ\phi}C\to{}_{\psi\circ\phi}C\boxtimes C$ is given by $(\mathbbm{w}*\mathbbm{v})_{1_{\mathcal{C}}}(a\boxtimes c)=\psi(\phi(a))\boxtimes c.$ By definition,
$(\mathbbm{w}*\mathbbm{v})_{1_{\mathcal{C}}}=(T\boxtimes\mathrm{id}_{C})\circ(\mathrm{id}_{B}\boxtimes\mathbbm{w}_{1_{\mathcal{C}}})\circ(\mathbbm{v}_{1_{\mathcal{C}}}\boxtimes\mathrm{id}_{C})\circ S_{1_{\mathcal{C}}}.$
Let $\eta_{\lambda}$ be an approximate unit for $B$. Then, a direct calculation shows that
$(\mathbbm{w}*\mathbbm{v})_{1_{\mathcal{C}}}(a\boxtimes c)$ $=\lim\limits_{\lambda}(T\boxtimes\mathrm{id}_{C})\circ(\mathrm{id}_{B}\boxtimes\mathbbm{w}_{1_{\mathcal{C}}})\circ(\mathbbm{v}_{1_{\mathcal{C}}}\boxtimes\mathrm{id}_{C})(a\boxtimes\eta_{\lambda}\boxtimes c)$
$=\lim\limits_{\lambda}(T\boxtimes\mathrm{id}_{C})\circ(\mathrm{id}_{B}\boxtimes\mathbbm{w}_{1_{\mathcal{C}}})(\phi(a)\boxtimes\eta_{\lambda}\boxtimes c)$
$=\lim\limits_{\lambda}(T\boxtimes\mathrm{id}_{C})(\phi(a)\boxtimes\psi(\eta_{\lambda})\boxtimes c)$
$=\lim\limits_{\lambda}\psi(\phi(a))\psi(\eta_{\lambda})\boxtimes c$
$=\psi(\phi(a))\boxtimes c,$
for any $a\in A$ and any $c\in C$. Hence, $(\psi\circ\phi,\mathbbm{w}*\mathbbm{v})$ is a cocycle morphism. ∎
@@
With the same notation as in Definition 3.6, if $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are extendible cocycle representations, then it also follows that $(\psi\circ\phi,\mathbbm{w}*\mathbbm{v})$ is a well-defined cocycle representation.
@@
If the cocycle morphisms in Definition 3.6 are assumed to be non-degenerate, then the maps $S_{X}$ and $T$ are bimodule isomorphisms for any $X\in\mathcal{C}$. Therefore, the composition in Definition 3.6 (denoted ***) corresponds canonically to the composition of correspondence morphisms (denoted $\circ$).
The composition considered in Definition 3.6 defines a category. To show this we first reformulate the notion of cocycle morphism. Roughly speaking, all the information carried by the collection of isometries $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ can be encoded into a collection of linear maps $\{h^{X}:F(X)\to G(X)\}_{X\in\mathcal{C}}$ satisfying some conditions. This viewpoint will facilitate our constructions and proofs in later sections. Our approach is motivated by [6, Lemma 3.8] which introduces this alternative viewpoint in the unital setting.
First, we recall a way of extending an action $(F,J)$ of a $*{}^{*}$-tensor category $\mathcal{C}$ on $A$ to its matrix amplification $M_{n}(A)$. Consider the functor $F^{\mathrm{amp}(n)}:\mathcal{C}\rightarrow\mathrm{Corr}(M_{n}(A))$ that maps objects $X\in\mathcal{C}$ to the correspondence with underlying bimodule $F(X)\otimes M_{n}(\mathbb{C})$ with the right inner product defined by $\langle(x_{ij}),(y_{ij})\rangle_{M_{n}(A)}=(\sum_{l}\langle x_{li},y_{lj}\rangle_{A})$, right $M_{n}(A)$ action given by $(x_{ij})\lhd(a_{ij})=(\sum_{l}x_{il}\lhd a_{lj})$ and with the left action given by $(a_{ij})\rhd(x_{ij})=(\sum_{l}a_{il}\rhd x_{lj})$ for $a_{ij}\in A$ and $x_{ij},y_{ij}\in F(X)$ for $1\leq i,j\leq n$ [^11] This bimodule can be identified with the external tensor product of $F(X)$ and $M_{n}(\mathbb{C})$. Hence the inner product defines a Hilbert $M_{n}(A)$-bimodule (see [28]). For a morphism $T\in\mathrm{Hom}(X,Y)$ we let $F^{\mathrm{amp}(n)}(T)=T\otimes\mathrm{id}_{M_{n}(\mathbb{C})}$. Moreover, letting $J_{X,Y}^{\mathrm{amp}(n)}((y_{ij})\boxtimes(x_{ij})))=(\sum_{l}J_{X,Y}(y_{il}\boxtimes x_{lj}))$ for $x_{ij}\in F(X)$ and $y_{ij}\in F(Y)$ for $1\leq i,j\leq n$, it is a straightforward calculation that $(F^{\mathrm{amp}(n)},J^{\mathrm{amp}(n)})$ is an action of $\mathcal{C}$ on $M_{n}(A)$.
@@
Let $\mathcal{C}$ be a $\mathrm{C}^{*}$-tensor category acting on $\mathrm{C}^{*}$-algebras $A$ and $B$ via $(A,F,J)$ and $(B,G,I)$ respectively, and let $\phi:A\rightarrow B$ be a $*{}^{*}$-homomorphism. Then there is a bijection between the families $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ corresponding to a cocycle morphism $(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ and families of linear maps:
@@
[…]
@@
Suppose we are given a collection of linear maps $\{h^{X}\}$ satisfying the conditions listed above. Fix $\zeta_{\lambda}$ an approximate unit for $A$. For all $X\in\mathcal{C}$ let $\mathbbm{v}_{X}:F(X)\boxtimes{}_{\phi}B\to{}_{\phi}B\boxtimes G(X)$ be given by
@@
Since the bimodule $F(X)$ is non-degenerate, denoting the left action by $\sigma:A\to\mathcal{L}(F(X))$, we have that $\sigma(\zeta_{\lambda})$ converges strictly to $1_{\mathcal{L}(F(X))}$. Then, for any $x\in F(X)$, $\zeta_{\lambda}\rhd x=\sigma(\zeta_{\lambda})x$ converges to $x$. Moreover, condition (iii) implies that $h^{X}$ is continuous, so the right hand side of (3.11) converges to zero and hence $\|\phi(\zeta_{\lambda}-\zeta_{\mu})\boxtimes h^{X}(x)\lhd b\|^{2}$ converges to 00. Therefore the formula in (3.10) gives a well-defined map $\mathbbm{v}_{X}$ for all $X\in\mathcal{C}$.
@@
Therefore, $\mathbbm{v}_{X}$ is an $A$-$B$-bimodule map. Moreover, using (v),
@@
(3.12) {tikzcd}
@@
Now, consider the map $\Psi:\{h^{X}\}\to\{\mathbbm{v}_{X}\}$ given by the formula in (3.10). We claim that $\Psi$ is independent of the choice of approximate unit. Indeed let $\zeta_{\lambda}$ and $\xi_{\lambda}$ be two approximate units for $A$. Similarly as in (3.11) we have that, $\|\phi(\zeta_{\lambda}-\xi_{\lambda})\boxtimes h^{X}(x)\lhd b\|=\|h^{X}(\zeta_{\lambda}\rhd x-\xi_{\lambda}\rhd x)\lhd b\|$, which converges to 00. Hence, the map $\mathbbm{v}_{X}$ is independent of the choice of approximate unit, and so $\Psi$ is well-defined.

### 08 Apr 2026 v3

@@ 3. The generalised cocycle category
In this section we introduce the category of $\mathcal{C}$-$\mathrm{C}^{*}$-algebras for which we will later perform intertwining arguments. Throughout the rest of this paper $\mathcal{C}$ is always assumed to be a $\mathrm{C}^{*}$-tensor category.
@@
Let $\mathcal{C}\overset{F}{\curvearrowright}A$ and $\mathcal{C}\overset{G}{\curvearrowright}B$ be actions of $\mathcal{C}$ on $\mathrm{C}^{*}$-algebras AA and $B$.
@@
    A correspondence morphism from $(A,F,J)$ to $(B,G,I)$ is a pair $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}})$, where $\phi:A\to\mathcal{L}(E)$ is an AA-$B$-correspondence and $\{\mathbbm{v}_{X}:\phi\circ F(X)\to G(X)\circ\phi\}_{X\in\mathcal{C}}$ is a natural family of AA-$B$-bimodule maps such that $\mathbbm{v}_{X}$ is an isometry (not necessarily adjointable) for any $X\in\mathcal{C}$. [^6] By an isometry, we mean a map which preserves the norm. By the proof of the theorem in [30], it is equivalent to assume that it preserves the inner product. Moreover, for all $X,Y\in\mathcal{C}$ the following pentagon diagram commutes
    (3.1) $\phi\circ F(X)\circ F(Y)$$G(X)\circ\phi\circ F(Y)$$\phi\circ F(X\otimes Y)$$G(X)\circ G(Y)\circ\phi$$G(X\otimes Y)\circ\phi,$$\scriptstyle{J_{X,Y}\boxtimes\mathrm{id}_{\phi}}$$\scriptstyle{\mathrm{id}_{F(Y)}\boxtimes\mathbbm{v}_{X}}$$\scriptstyle{\mathbbm{v}_{Y}\boxtimes\mathrm{id}_{G(X)}}$$\scriptstyle{\mathbbm{v}_{X\otimes Y}}$$\scriptstyle{\mathrm{id}_{\phi}\boxtimes I_{X,Y}}$
@@
    A cocycle representation $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ is a correspondence morphism for which we further require that $\phi:A\to\mathcal{M}(B)$ is a -homomorphism.
@@
    A cocycle morphism $(\phi,\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}):(A,F,J)\to(B,G,I)$ is a cocycle representation for which we further require that $\phi:A\to B$ is a -homomorphism.
@@
Example $3.$
Suppose $(A,\alpha)$ and $(B,\beta)$ are actions of a countable discrete group $\Gamma$ on $\mathrm{C}^{*}$-algebras AA and $B$. Consider them as actions of $\mathrm{Hilb}(\Gamma)$ as in Example 2.4. Let $(\phi,\mathbbm{v}):(A,\alpha)\to(B,\beta)$ be an extendible cocycle morphism as in Definition 3.1 with $\mathbbm{v}_{\mathbb{C}g}$ being adjointable for all $g\in\Gamma$. Fix $g\in\Gamma$ and let $f_{g}:=T_{g}^{\prime}\circ\mathbbm{v}_{\mathbb{C}_{g}}\circ S_{g}:{}_{\phi\circ\alpha_{g}}B\to{}_{\beta_{g}\circ\phi}B$, where $S_{g}:{}_{\phi\circ\alpha_{g}}B\to{}_{\alpha_{g}}A\boxtimes{}_{\phi}B$ is given by $S_{g}(b)=\lim\limits_{\lambda}\xi_{\lambda}\boxtimes b$ for any $b\in B$, where $\xi_{\lambda}$ is an approximate unit for AA and $T_{g}^{\prime}:{}_{\phi}B\boxtimes{}_{\beta_{g}}B\rightarrow{}_{\beta_{g}\circ\phi}B$ is given by $T_{g}^{\prime}(b\boxtimes c)=\beta_{g}(b)c$ for $b,c\in B$. [^8] Note that $S_{g}$ is well-defined by Proposition 1.6. Moreover, let $T_{g}:{}_{\alpha_{g}}A\boxtimes{}_{\phi}B\rightarrow{}_{\phi\circ\alpha_{g}}B$ be given by $T_{g}(a\boxtimes b)=\phi(a)b$ for any $a\in A$ and any $b\in B$.
@@
(3.2) $_{\alpha_{g}}A\boxtimes{}_{\phi}B$$_{\phi}B\boxtimes{}_{\beta_{g}}B$$_{\phi\circ\alpha_{g}}B$$_{\beta_{g}\circ\phi}B$$\scriptstyle{T_{g}}$$\scriptstyle{\mathbbm{v}_{\mathbb{C}_{g}}}$$\scriptstyle{f_{g}}$$\scriptstyle{(T_{g}^{\prime})^{-1}}$
@@
Note that, unlike in [41, Definition 1.10], we do not require $\phi\in\mathrm{Corr}(A,B)$ to be extendible. In fact, this is precisely the reason why we consider the maps $\mathbbm{v}_{X}$ to be isometries (possibly non-adjointable) instead of unitaries. For example, the map $\mathbbm{v}_{1_{\mathcal{C}}}:A\boxtimes{}_{\phi}B\to{}_{\phi}B\boxtimes B\cong{}_{\phi}B$ is given by $\mathbbm{v}_{1_{\mathcal{C}}}(a\boxtimes b)=\phi(a)b$ for all $a\in A$ and $b\in B$. Therefore, it is not surjective unless $\phi$ is non-degenerate. Moreover, it might not be adjointable if $\phi$ is not extendible as seen in the proof of Proposition 1.6.
Furthermore, the morphisms of Definition 3.1 fit into the $\mathcal{C}$-equivariant $\mathrm{KK}$-theory developed in [1] (see [1, Example 3.3]). In [1] a correspondence morphism is instead called a $\mathcal{C}$-Hilbert AA-$B$-bimodule and a cocycle morphism is called a cocycle-$\mathcal{C}$-*-homomorphism.
We now define composition formulae for the various notions of morphisms in Definition 3.1. Using the standard composition between correspondences, we can define composition of correspondence morphisms in the obvious way. Let $\mathcal{C}\overset{F}{\curvearrowright}A$, $\mathcal{C}\overset{G}{\curvearrowright}B$, and $\mathcal{C}\overset{H}{\curvearrowright}C$ be actions of $\mathcal{C}$ on $\mathrm{C}^{*}$-algebras AA, $B$, and $C$ respectively. If $(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ and $(\psi,\mathbbm{w}):(B,G,I)\to(C,H,K)$ are correspondence morphisms, their composition is denoted by $(\psi\circ\phi,\mathbbm{w}\circ\mathbbm{v})$, where
@@
By combining the pentagon diagrams for $\mathbbm{v}$ and $\mathbbm{w}$, one obtains that $(\psi\circ\phi,\mathbbm{w}\circ\mathbbm{v}):(A,F,J)\to(C,H,K)$ is indeed a correspondence morphism.
@@
(3.4) $F(X)\boxtimes{}_{\psi\circ\phi}C$$_{\psi\circ\phi}C\boxtimes H(X)$$F(X)\boxtimes{}_{\phi}B\boxtimes{}_{\psi}C$$_{\phi}B\boxtimes{}_{\psi}C\boxtimes H(X).$$\scriptstyle{S_{X}}$$\scriptstyle{(\mathbbm{w}*\mathbbm{v})_{X}}$$\scriptstyle{(\mathbbm{w}\circ\mathbbm{v})_{X}}$$\scriptstyle{T\boxtimes\mathrm{id}_{H(X)}}$
Here $S_{X}(x\boxtimes c)=\lim\limits_{\lambda}x\boxtimes\eta_{\lambda}\boxtimes c$ and $T(b\boxtimes c)=\psi(b)c$ for any $X\in\mathcal{C}$, $x\in F(X)$, $b\in B$, and $c\in C$, with $\eta_{\lambda}$ being an approximate unit for BB. [^10] We have not shown that $S_{X}$ is well-defined at this point.
@@
By the Cauchy-Schwarz inequality and since $(\eta_{\lambda}-\eta_{\mu})\phi(\langle x,x\rangle_{A})^{1/2}$ converges to 0, it is readily seen that $\langle x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c,x\boxtimes(\eta_{\lambda}-\eta_{\mu})\boxtimes c\rangle_{C}$ converges to 0, so $S_{X}$ is well-defined. Moreover, commutation with the left and right actions are immediate.
@@
Therefore, by linearity and continuity of both $S_{X}$ and $T_{X}$, it follows that $S_{X}$ is surjective and $T_{X}$ is injective. It now suffices to show that $T_{X}$ is surjective. This will imply that $S_{X}$ is invertible with $T_{X}$ being the inverse. Let $x\in F(X)$ and $c\in C$. Since $F(X)$ is a non-degenerate AA-bimodule, the map $R_{X}^{-1}:F(X)\boxtimes A\to F(X)$ given by $R_{X}^{-1}(x\boxtimes a)=x\lhd a$ is a bimodule isomorphism (see the discussion above Lemma 1.10). Then, let $y\in F(X)$ and $a\in A$ such that $y\lhd a=x$. A straightforward calculation shows that $T_{X}(y\boxtimes\phi(a)\boxtimes c)=x\boxtimes c$, so $T_{X}$ is surjective. Hence, $S_{X}=T_{X}^{-1}$ is an isomorphism for any $X\in\mathcal{C}$.
@@
Note that the proof of Lemma 3.6 also shows that the map $S_{X}$ does not depend on the choice of approximate unit. The proof of the following Lemma is routine.
@@
We invite the reader to recall Remark 1.5. Then, with the same notation as in Definition 3.5, if $(\phi,\mathbbm{v})$ and $(\psi,\mathbbm{w})$ are cocycle representations and $\psi$ is extendible, then it also follows that $(\psi\circ\phi,\mathbbm{w}*\mathbbm{v})$ is a well-defined cocycle representation.
@@
If the cocycle morphisms in Definition 3.5 are assumed to be non-degenerate, then the maps $S_{X}$ and $T$ are bimodule isomorphisms for any $X\in\mathcal{C}$. Therefore, the composition in Definition 3.5 (denoted *) corresponds canonically to the composition of correspondence morphisms (denoted $\circ$).
The composition considered in Definition 3.5 defines a category. To show this we first reformulate the notion of cocycle morphism. Roughly speaking, all the information carried by the collection of isometries $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ can be encoded into a collection of linear maps $\{h^{X}:F(X)\to G(X)\}_{X\in\mathcal{C}}$ satisfying some conditions. This viewpoint will facilitate our constructions and proofs in later sections. Our approach is motivated by [7, Lemma 3.8] which introduces this alternative viewpoint in the unital setting.
First, we recall a way of extending an action $(F,J)$ of a C-tensor category $\mathcal{C}$ on AA to its matrix amplification $M_{n}(A)$. Consider the functor $F^{\mathrm{amp}(n)}:\mathcal{C}^{\mathrm{rev}}\rightarrow\mathrm{Corr}(M_{n}(A))$ that maps objects $X\in\mathcal{C}$ to the correspondence with underlying bimodule $F(X)\otimes M_{n}(\mathbb{C})$ with the right inner product defined by $\langle(x_{ij}),(y_{ij})\rangle_{M_{n}(A)}=(\sum_{l}\langle x_{li},y_{lj}\rangle_{A})$, right $M_{n}(A)$ action given by $(x_{ij})\lhd(a_{ij})=(\sum_{l}x_{il}\lhd a_{lj})$ and with the left action given by $(a_{ij})\rhd(x_{ij})=(\sum_{l}a_{il}\rhd x_{lj})$ for $a_{ij}\in A$ and $x_{ij},y_{ij}\in F(X)$ for $\leq i,j\leq n.11$1111This bimodule can be identified with the external tensor product of $F(X)$ and $M_{n}(\mathbb{C})$. Hence the inner product defines a Hilbert $M_{n}(A)$-bimodule (see [29]). For a morphism $T\in\mathrm{Hom}(X,Y)$ we let $F^{\mathrm{amp}(n)}(T)=T\otimes\mathrm{id}_{M_{n}(\mathbb{C})}$. Moreover, letting $J_{X,Y}^{\mathrm{amp}(n)}((y_{ij})\boxtimes(x_{ij}))=(\sum_{l}J_{X,Y}(y_{il}\boxtimes x_{lj}))$ for $x_{ij}\in F(X)$ and $y_{ij}\in F(Y)$ for $1\leq i,j\leq n$, it is a straightforward calculation that $(F^{\mathrm{amp}(n)},J^{\mathrm{amp}(n)})$ is an action of $\mathcal{C}$ on $M_{n}(A)$.
@@
Let $\mathcal{C}$ be a $\mathrm{C}^{*}$-tensor category acting on $\mathrm{C}^{*}$-algebras AA and $B$ via $(A,F,J)$ and $(B,G,I)$ respectively, and let $\phi:A\rightarrow B$ be a -homomorphism. Then there is a bijection between the families $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ corresponding to a cocycle morphism $(\phi,\mathbbm{v}):(A,F,J)\to(B,G,I)$ and families of linear maps:
@@
    $F(Y)\boxtimes F(X)$$F(X\otimes Y)$$G(Y)\boxtimes G(X)$$G(X\otimes Y)$$\scriptstyle{h^{Y}\boxtimes h^{X}}$$\scriptstyle{J_{X,Y}}$$\scriptstyle{h^{X\otimes Y}}$$\scriptstyle{I_{X,Y}}$
@@
Suppose we are given a collection of linear maps $\{h^{X}\}$ satisfying the conditions listed above. Fix $\zeta_{\lambda}$ an approximate unit for AA. For all $X\in\mathcal{C}$ let $\mathbbm{v}_{X}:F(X)\boxtimes{}_{\phi}B\to{}_{\phi}B\boxtimes G(X)$ be given by
@@
Since the bimodule $F(X)$ is non-degenerate, denoting the left action by $\sigma:A\to\mathcal{L}(F(X))$, we have that $\sigma(\zeta_{\lambda})$ converges strictly to $1_{\mathcal{L}(F(X))}$. Then, for any $x\in F(X)$, $\zeta_{\lambda}\rhd x=\sigma(\zeta_{\lambda})x$ converges to $x$. Moreover, condition (iii) implies that $h^{X}$ is continuous, so the right hand side of (3.6) converges to zero and hence $\|\phi(\zeta_{\lambda}-\zeta_{\mu})\boxtimes h^{X}(x)\lhd b\|^{2}$ converges to 0. Therefore the formula in (3.5) gives a well-defined map $\mathbbm{v}_{X}$ for all $X\in\mathcal{C}$.
@@
Therefore, $\mathbbm{v}_{X}$ is an AA-$B$-bimodule map. Moreover, using (v),
@@
(3.7) $F(Y)\boxtimes F(X)\boxtimes{}_{\phi}B$$F(Y)\boxtimes{}_{\phi}B\boxtimes G(X)$$F(X\otimes Y)\boxtimes{}_{\phi}B$$_{\phi}B\boxtimes G(Y)\boxtimes G(X)$$_{\phi}B\boxtimes G(X\otimes Y).$$\scriptstyle{J_{X,Y}\boxtimes\mathrm{id}_{B}}$$\scriptstyle{\mathrm{id}_{F(Y)}\boxtimes\mathbbm{v}_{X}}$$\scriptstyle{\mathbbm{v}_{Y}\boxtimes\mathrm{id}_{G(X)}}$$\scriptstyle{\mathbbm{v}_{X\otimes Y}}$$\scriptstyle{\mathrm{id}_{B}\boxtimes I_{X,Y}}$
@@
Now, consider the map $\Psi:\{h^{X}\}\to\{\mathbbm{v}_{X}\}$ given by the formula in (3.5). We claim that $\Psi$ is independent of the choice of approximate unit. Indeed let $\zeta_{\lambda}$ and $\xi_{\lambda}$ be two approximate units for AA. Similarly as in (3.6) we have that, $\|\phi(\zeta_{\lambda}-\xi_{\lambda})\boxtimes h^{X}(x)\lhd b\|=\|h^{X}(\zeta_{\lambda}\rhd x-\xi_{\lambda}\rhd x)\lhd b\|$, which converges to 0. Hence, the map $\mathbbm{v}_{X}$ is independent of the choice of approximate unit, and so $\Psi$ is well-defined.